OpenStudy (anonymous):

In which quadrant will the point of intersection lie? 2x-y=3 and x-2y=5

OpenStudy (tkhunny):

Intercept Form should give it away... A: \(\dfrac{x}{3/2} + \dfrac{y}{-3} = 1\) B: \(\dfrac{x}{5} + \dfrac{y}{-2/5} = 1\) y-intercept A(-3) is farther from the Origin than B(-2/5) x-intercept A(3/2) is closer to the Origin than B(5) They must have crossed in Quadrant IV.